# Multi-modal Retrieval Augmented Multi-modal Generation: A Benchmark, Evaluate Metrics and Strong Baselines

### v1

@@ 8 Analysis
In this section, we answer the following three important questions about our proposed approaches.
8.1 Topic is a Key Factor Affecting Performance
The average overall scores of all evaluated models, as presented in Figure 3, reveal notable variations in performance across different topics. For instance, the most challenging topic is “Business & Finance”, which achieves the lowest overall score, whereas “Society & Culture” and “Sports” are significantly easier topics by comparison.
@@ 8.1 Topic is a Key Factor Affecting Performance
We hypothesize that these performance differences stem from the varying demands for multi-modal information across topics. Specifically, images are more prevalent and useful in less technical fields, such as sports and culture. In contrast, fields like business and finance often require data from additional modalities, such as tables and charts, which are less readily integrated. Moreover, theoretical fields like politics may not rely heavily on multi-modal information for comprehensive responses.
To achieve balanced performance across topics, a topic-sensitive approach should be developed to adaptively incorporate appropriate modalities for response synthesis, ensuring that the specific requirements of each topic are effectively addressed.
8.2 Caption Quality is Important for Image Integration
As described in Section 5.2, images are converted into captions for use by LLMs. To ensure a comparable level of information between LLMs and MLLMs, the captions are crafted to be as detailed as possible. Additionally, textual contexts associated with the images are included during the captioning process to enhance consistency. In this section, we investigate the impact of removing these elements.
Figure 4: Comparison of performance with original and ablation settings.
As shown in Figure 4, performance declines noticeably with the removal of details and contextual information, particularly in the metrics of Image Coherence and Image Reference. The absence of contextual information in captions leads to a significant drop in Image Recall, highlighting the critical role of context in guiding the image selection process of LLMs. Beyond these key observations, the ablation settings show minor fluctuations in other metrics. Overall, incorporating context during captioning is essential for achieving optimal image retrieval, while including detailed descriptions further enhances performance to a measurable extent.
8.3 Contributions of Auxiliary Images
Auxiliary images are collected for each query to address cases where highly relevant images are absent from the retrieved web pages. Removing these images leads to a significant decline in generative performance (Figure 4). For instance, Image Recall drops sharply without auxiliary images, suggesting that LLMs preferentially include auxiliary images over those sourced from web pages when constructing the generated document. Additionally, the decline in Image Coherence and Image Helpfulness metrics highlights the superior quality of auxiliary images compared to web-sourced images. Interestingly, the Image Reference metric shows a notable increase, likely due to stronger connections between each remaining image and its context. Overall, the inclusion of auxiliary images significantly enriches the visual content and enhances the quality of the generated document.
@@ 10 Limitations
Topic Restrictions
Certain topics, such as “Business & Finance” or “Politics & Government”, are not well-suited for visual representation. For these domains, tables, charts, or other formats may be more appropriate than images. The absence of these modalities limits the applicability of our approach to these areas.
Content Availability
The lack of extensive public discussion or data in certain specialized fields presents challenges for retrieving abundant multi-modal documents effectively. This limitation reduces the comprehensiveness of our analysis in underrepresented domains.
@@
The proposed evaluation metrics are limited and may fail to capture all aspects of M2RAG evaluation. For example, pairwise ranking is more appropriate for assessing overall performance than Likert scale in M2RAG task, which lacks a standard optimal answer. However, using the LLM-as-a-Judge Zheng et al. (2023) approach, pairwise comparisons are resource-intensive, particularly for large-scale evaluations. Furthermore, their high cost and impracticality make them unsuitable for benchmarking, necessitating reliance on fixed scoring metrics that may not adequately address all edge cases.
Appendix A Query Collection
@@ 1. Society & Culture
Appendix B Retrieval of Collected Elements
We replaced traditional embedding-based retrieval methods with LLM-based and MLLM-based approaches to enhance the retrieval process. Embedding-based methods typically prioritize semantic similarity but often overlook whether candidate elements provide substantial information relevant to the target question. Additionally, in real-time M2RAG tasks, where online documents cannot be pre-indexed, LLM-based and MLLM-based methods offer better performance without incurring significant additional time costs.
Using these methods, text and visual elements within the collected documents are evaluated based on their relevance to the user query. This evaluation is conducted by prompting LLMs and MLLMs for textual and visual content, respectively. The prompt templates used for these evaluations are illustrated in Figures 8 and 9.
@@ 2. Text from a Web Page # Evaluation Guidelines
Appendix C Dataset Statistics
C.1 Basic Statistics
@@
Item Value
Queries 200
Web pages per query 6.4
Text elements per web page 7.2
Images per query 9.3
Auxiliary image per query 2.6
Average score of text elements 1.1
Average score of image elements 4.2
Average auxiliary image score 3.6
@@ C.2 Data Distribution
Appendix D Details of Evaluation Metrics
@@
You are a multi-modal document evaluation assistant. You will receive an image and its textual context written in markdown. Your task is to evaluate the coherence between the image and the text (context above and below) it accompanies.
@@
Assess how coherent the image is in relation to its accompanying text, assigning a score from 0 to 10. A higher score indicates stronger coherence between the image and the text. Be precise when assigning the score.
- A score from 0-3 means that the image is minimally or not at all coherent with the text. - A score from 4-6 indicates that the image shows some coherence with the text but may include unrelated elements. - A score from 7-9 indicates that the image is highly coherent with the text. - A score of 10 indicates perfect coherence, where the image completely corresponds with and enhances the text. Be rigorous and discerning when assigning your score. # Output Instructions
Your output should contain only 2 lines:
Figure 19: Prompt template for the image coherence evaluation.
# Task Description
You are a multi-modal document evaluation assistant. You will receive an image and its textual context written in markdown. Your task is to evaluate the helpfulness of the image in enabling human readers to comprehend the text (context above and below) it accompanies.
# Context Above
markdown
{context_above} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Context Below
markdown
{context_below} ‘‘‘ # Scoring Criteria
Evaluate how well the image helps human readers understand the content of its accompanying text, assigning a score from 0 to 10. A higher score indicates that the image significantly enhances comprehension of the text. Be as precise as possible when assigning the score.
- 0-3: The image is minimally or not at all helpful for comprehension. - 4-6: The image provides some helpful context or information but may contain extraneous or less relevant details. - 7-9: The image is highly helpful in enabling comprehension of the text. - 10: The image is perfectly helpful and greatly enhances and clarifies the information provided in the text. Be rigorous and discerning when assigning your score. # Output Instructions
Your output should contain only 2 lines:
Figure 20: Prompt template for the image helpfulness evaluation.
# Task Description
You are a multi-modal document quality assessment assistant. You will receive an image and its accompanying textual context, formatted in markdown. Your task is to determine whether the image is explicitly referenced or explained within the surrounding text (both above and below the image).
# Context Above
markdown
{context_above} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Context Below
markdown
{context_below} ‘‘‘ # Scoring Criteria
Evaluate the extent to which the image is referenced or explained in the text:
- 0: The image is not mentioned or referenced in the context. - 1-3: The image is referenced implicitly, and the reference is improper or incorrect. - 4-6: The image is referenced explicitly but in an improper manner, or it is referenced implicitly. - 7-9: The image is referenced explicitly, with the reference being generally proper and correct. - 10: The image is referenced explicitly, with the placement and explanation being completely proper and correct. Be rigorous and discerning when assigning your score. # Output Instructions
@@
# Task Description
You are a multi-modal document evaluation assistant. You will receive a single image and a multi-modal document written in markdown, where detailed captions serve as proxies for images. Your task is to evaluate whether the image should be inserted into the document to enhance understanding for human readers.
# Images in the Document
The images in the document have been replaced with their detailed captions. Each image is represented by a single line starting with ‘<image>:’.
# Document
markdown
{text} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Scoring Criteria
Evaluate whether the provided image should be inserted into the document based on the following:
- If the image is irrelevant to the document’s content or does not aid in comprehension, output a single integer 0, indicating the image should not be inserted. - If the image content is too similar to or repetitive of images already represented in the document, output a single integer 0, indicating the image should not be inserted. - Otherwise, output a single integer 1, indicating the image should be inserted. # Output Instructions
Your output should contain only two lines:
1. A brief explanation justifying the score of 0 or 1. 2. The score as an integer value. Do not provide any additional information beyond the explanation and score.
Figure 22: Prompt template for the image recall evaluation.
Appendix E Experimental Setup

### v2

@@ 6 Analysis
In this section, we present more analysis on our benchmark by answering several questions.
6.1 Does Topic Affect Performance?
The average overall scores of all evaluated models, as presented in Figure 3, reveal notable variations in performance across different topics. For instance, the most challenging topic is Politics & Government, which achieves the lowest overall score, whereas Society & Culture and Sports are significantly easier topics by comparison.
@@ 6.1 Does Topic Affect Performance?
We hypothesize that these performance differences stem from the varying demands for multi-modal information across topics. Specifically, images are more prevalent and useful in less technical fields, such as sports and culture. In contrast, fields like politics and finance often require more accurate data from additional modalities, such as tables and charts, which are less readily integrated. These phenomena suggest that future work needs to carefully implement strategies to control the proportion of included image across different topics.
6.2 Is Description Quality Important?
As described in Section 4.3, to provide the LLMs with rich and accurate image content, each image is transcribed into a detailed textual description, with consideration of its original contextual information. In this subsection, we conduct ablation studies to investigate the contributions of these information. Table 3 reveals that model’s performance significantly declines by removing contextual information, particularly in the text-modal and multi-modal precision metrics. The absence of details in descriptions leads to a notable drop in Image Recall, highlighting the critical role of details in guiding image selection of LLMs. Overall, incorporating context is essential for achieving optimal image retrieval, while including detailed descriptions further enhances performance to a measurable extent.
Ablation
Text
Avg.
Multi-Modal
Precision Avg.
Recall
Score
Overall
Score
Original 88.9 56.1 75.8 74.9
w/o Context 88.5 51.4 76.8 73.1
w/o Detail 88.8 54.4 73.5 74.0
w/o Aux. Image 88.7 52.4 - -
Table 3: Comparison of GPT-4o experiment and ablation study results. Text Avg. and Multi-Modal Precision Avg. represent the average values of four text-modal metrics and three multi-modal precision metrics (except for Recall), respectively. Recall and Overall scores for w/o Aux. Image is not reported due to changes in the images within the knowledge base.
6.3 Are Auxiliary Images Helpful?
Auxiliary images are collected for each query to address cases where highly relevant images are absent from the retrieved web pages. As illustrated in last row of Table 3 (w/o Aux. Image), removing these auxiliary images leads to a significant decline in generative performance in multi-modal metrics. This observation reflects the challenges in effectively integrating auxiliary images that lack contextual information with the textual content. Overall, the inclusion of auxiliary images significantly enriches the visual content and enhances the quality of the generated document.
@@ 7 Conclusion
More Modalities
Our work mainly focus on the text and image modalities. In the real-world scenarios, information in other modalities are also crucial, such as speech and video. In the future, we plan to extend our proposed benchmark and training datasets to more modalities, even the mixed modalities.
@@
The proposed evaluation metrics might fail to capture all potential aspects of M2RAG evaluation. For example, pairwise ranking is more appropriate for assessing overall performance than Likert scale in M2RAG task, which lacks a standard optimal answer. Besides, due to the requirement of using LLMs and MLLMs to evaluate the quality of long-text multi-modal content across 8 dimensions [^7] Please refer to Table 6 for more details about evaluation cost., this leads to huge evaluation costs in large-scale evaluation. To balance the affordability and reliability of experiments on our proposed benchmark, we introduce a quantity-first approach with 750 evaluation samples in this paper. In future work, we plan to distill and build efficient text-modal and multi-modal metrics to support larger-scale benchmark testing.
Appendix A Ethical Considerations
Most of the task inputs in our benchmark and training dataset are sourced from publicly available datasets, ensuring that they pose no harm to individuals or groups. Furthermore, the text generated by large language models (LLMs) and multi-modal language models (MLLMs) is carefully curated and processed by human annotators to safeguard privacy and confidentiality. No personally identifiable information (PII) is included. However, it is important to note that texts from the ELI5 dataset Fan et al. (2019) and multi-modal documents retrieved via Google Custom Search may contain harmful content or hate speech. Despite these potential risks, it is crucial to disclose the full scope of this research, as materials from ELI5 and Google Custom Search have been extensively used in safety research within the community. All annotators are compensated fairly, with an hourly wage of approximately \$4.76 USD, which exceeds the average hourly wage of \$3.13 USD on Amazon Mechanical Turk Hara et al. (2018).
Appendix B Query Collection
@@ 1. Society & Culture
ID Category Name Num. Samples
1 Society & Culture 14182
2 Science & Mathematics 14449
3 Health 10931
4 Education & Reference 776
5 Computers & Internet 6999
6 Sports 1974
7 Business & Finance 3821
8 Entertainment & Music 1946
9 Family & Relationships 820
10 Politics & Government 4638
11 Others 451
Table 4: Topics and volumes of the queries.
Appendix C Retrieval of Collected Elements
We replaced traditional embedding-based retrieval methods with LLM-based and MLLM-based approaches to improve the retrieval performance. Embedding-based methods typically prioritize semantic similarity but often overlook whether candidate elements provide substantial information relevant to the target question. Additionally, in real-time M2RAG tasks, where online documents cannot be pre-indexed, LLM-based and MLLM-based methods offer better performance without incurring significant additional time costs.
Using these methods, text and visual elements within the collected documents are evaluated based on their relevance to the user query. This evaluation is conducted by prompting LLMs and MLLMs for textual and visual content, respectively. The prompt templates used for these evaluations are illustrated in Figures 7 and 8. This step results in a score from 0 to 10 for each element, where higher scores indicate better relevancy.
@@ 2. Text from a Web Page # Evaluation Guidelines
Appendix D Dataset Statistics
D.1 Basic Statistics
@@
Item Range Value
Query benchmark 750
training 1605
Web Page per query 9.8
Image 12.0
Aux. Image 3.0
Text Element per web page 25.7
Image 1.2
Avg. Element Score text 1.5
image 3.1
aux. image 3.8
@@ D.2 Data Distribution
Appendix E Details of Evaluation Metrics
E.1 Prompt Templates
@@
You are a multimodal document evaluation assistant. You will receive an image and its textual context written in markdown. Your task is to evaluate the coherence between the image and the text (context above and below) it accompanies.
# Context Above
markdown
{context_above} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Context Below
markdown
{context_below} ‘‘‘ # Scoring Criteria and Examples
Evaluate how well the image complements its accompanying text and assign a score from 0 to 10. A higher score reflects stronger coherence between the image and the text. Be precise in your evaluation.
- Score 0-3: Minimal or no coherence between the image and text. - Example 1: The image shows a video game controller, but the text discusses supply and demand in virtual goods on platforms like Steam. These topics are unrelated, and the image adds no context to the discussion of virtual goods. The score should be 0-3. - Example 2: The image depicts a luxury car, but the text is about the environmental impact of electric vehicles. While both relate to automobiles, the image doesn’t contribute to the discussion of electric vehicles’ environmental effects. Thus, it merits a score of 0-3. - Example 3: A photo of a pizza is shown, but the text is a deep dive into the history of ancient Greece. There is no connection between the image and the historical content, making the coherence score 0-3. - Score 4-6: Some coherence, but with unrelated elements. - Example 1: The image shows various components like a remote control, cables, and a media player, which are related to the media player mentioned in the text. However, the context focuses on video formats, which is not addressed by the image. The coherence score should be between 4 and 6. - Example 2: The image shows graduates from Harvard University, and the text is about the cost of university education and the perceived benefits of attending prestigious institutions. While the image relates to the general topic of education, it doesn’t directly support the discussion about the financial costs or value of education, meriting a score of 4-6.
- Example 3: The image of a crowded subway is referenced in the text discussing urbanization and its effects on public transport. However, the text focuses more on the environmental impact of urbanization rather than directly linking to the image of subway crowds, leading to a coherence score of 4-6. - Score 7-9: High coherence, with the image closely aligning with the text. - Example 1: The image is a photo of a video game controller, and the text describes features of the latest gaming console. The image directly aligns with the content, visually representing the technology described in the text, earning a score of 7-9. - Example 2: The image depicts a close-up of a laptop keyboard, and the text discusses advancements in laptop design. The image complements the text well, helping the reader understand the specific design features mentioned, justifying a score of 7-9. - Example 3: A chart showing sales data for various smartphone models is paired with text explaining smartphone market trends. The image is directly relevant to the discussion of market share and trends, aiding comprehension, and warrants a score of 7-9. - Score 10: Perfect coherence, where the image completely corresponds to and enhances the text. - Example 1: The image is a detailed picture of a video game controller, and the text provides a comprehensive review of the controller’s features, design, and performance. The image fully complements the text, making the reader’s understanding more complete and vivid, deserving a score of 10. - Example 2: The image shows a close-up of a public toilet seat, which is relevant to the context below that discusses the design differences between public and home toilet seats. The image is completely aligned with the text, directly aiding the discussion. The score should be 10. - Example 3: The image is an anatomical diagram of the human heart, and the text explains how blood circulates through the body. The visual directly supports the explanation, making the text easier to understand and reinforcing key points, justifying a score of 10. Note: The image needs to be coherent with the context above OR below it, not necessarily both. Be rigorous and objective in your assessment. # Output Instructions
Your output should contain only 2 lines:
Figure 18: Prompt template for the image coherence evaluation.
# Task Description
You are a multimodal document evaluation assistant. You will receive an image and its textual context written in markdown. Your task is to evaluate the helpfulness of the image in enabling human readers to comprehend the text (context above and below) it accompanies.
# Context Above
markdown
{context_above} ‘‘‘ # Image
<IMAGE_PLACEHOLDER>
# Context Below
markdown
{context_below} ‘‘‘ # Scoring Criteria and Examples
Assess the image’s helpfulness in improving comprehension of the text, assigning a score from 0 to 10. A higher score indicates the image significantly enhances understanding. Be precise and rigorous when assigning the score.
- Score 0-3: The image is minimally or not at all helpful for comprehension. - Example 1: The image depicts a book cover titled "When God Asks You to Do Something You Don’t Want to Do," with a background of crashing ocean waves. The text is about jury members addressing trial observations. The image is irrelevant and unhelpful, meriting a score of 0-3. - Example 2: The image shows a close-up of banknotes (euro), arranged in a fan-like pattern. While the image is related to the concept of money, it does not directly enhance the comprehension of the text, which discusses the exchange rate of different currencies. Thus, it is not helpful for comprehension, earning a score of 0-3. - Example 3: The image is of a person looking at a blank computer screen, with no connection to the text discussing how the internet has influenced modern literature. The image does not provide any useful context or additional understanding of the text, so it warrants a score of 0-3. - Score 4-6: The image provides some helpful context but may include extraneous or less relevant details. - Example 1: The image shows a man at a desk, shocked at his laptop, with colorful text and symbols in the background. The text discusses how hackers are caught. While the image is thematically related, it is not directly helpful in explaining the topic of cyber security, earning a score of 4-6. - Example 2: The image shows a car with a vented hood, which relates to the text discussing the aerodynamic and cooling benefits of such designs in mid-engine cars. However, the image doesn’t address specific details, like material considerations or the pressure zones discussed in the text, meaning it provides partial context but not a complete understanding. Therefore, it should be assigned a score of 4-6.
- Example 3: The image of a group of college students at a campus event is referenced in the text discussing the benefits of a liberal arts education. While it offers a visual representation of student life, it doesn’t fully enhance the understanding of the specific academic benefits described, so it earns a score of 4-6. - Score 7-9: The image is highly helpful in enhancing comprehension of the text. - Example 1: The image features a table comparing popular browsers and their operating systems, such as Edge, Safari, and Chrome. The text discusses why Internet Explorer (IE) is disliked. The image is highly relevant, providing a clear visual context that directly supports the reader’s understanding of the browser landscape, justifying a score of 7-9. - Example 2: The image is a close-up of a laptop keyboard, with highlighted keys. The text discusses the evolution of laptop designs, and the image directly supports the explanation, helping the reader understand the physical changes over time, which justifies a score of 7-9. - Example 3: The image shows the front page of a scientific journal, with headlines and an image illustrating the main topic. The text explains the content of the journal in detail, and the image directly helps readers better visualize the article’s focus, justifying a score of 7-9. - Score 10: The image perfectly enhances and clarifies the text. - Example 1: The image is a detailed diagram of the human digestive system, showing organs like the stomach and intestines. The text explains the process of digestion, and the diagram complements this explanation perfectly, enhancing the reader’s understanding of the content. The image and text work together seamlessly, earning a score of 10. - Example 2: The image shows a close-up of a public toilet seat, which is relevant to the context below that discusses the design differences between public and home toilet seats. The image directly complements the text, providing the exact visual context needed to understand the design differences. The score should be 10. - Example 3: The image is a map highlighting the trade routes discussed in the text about ancient civilizations. The map directly enhances the comprehension of the trade dynamics and locations mentioned in the text, offering a visual aid that clarifies the complex subject matter, thus earning a score of 10. Note: The image needs to be helpful for the context above OR below it, not necessarily both. Be rigorous and objective in your assessment. # Output Instructions
Your output should contain only 2 lines:
Figure 19: Prompt template for the image helpfulness evaluation.
# Task Description
You are a multimodal document quality assessment assistant. You will receive an image and its accompanying textual context, formatted in markdown. Your task is to determine whether the image is explicitly referenced or explained within the surrounding text (both above and below the image).
@@
Determine how well the image is referenced or explained in the surrounding text, assigning a score from 0 to 10:
- Score 0: The image is not mentioned or referenced in the text. - Example 1: The image is present, but there is no mention or acknowledgment of it in either the context above or below. - Example 2: The image shows a scenic mountain view, but the text is entirely about office workflows, with no reference to or explanation of the image. - Example 3: The image is of a fruit basket, but the text is about urban planning, without any reference to the image or its context. - Score 1-3: The image is referenced implicitly, but the reference is inapparent, improper, or incorrect. - Example 1: The text discusses the Millennium Falcon’s hyperdrive system and kyber crystal energy field, which are depicted in the image. However, the image is not explicitly mentioned, and the connection is weak. The score is 1-3. - Example 2: The text mentions "the latest in gaming technology," and the image shows a futuristic gaming console. The image is not directly referenced, and the connection to the text is not clearly established. The score is 1-3. - Example 3: The text talks about space exploration and mentions rockets, but the image is a close-up of a satellite. The connection between the satellite and the discussion of space exploration is implied but not directly referenced. Thus, the score is 1-3. - Score 4-6: The image is referenced implicitly or explicitly, but the reference is improper or partially relevant.
- Example 1: The image shows a skier, and the text explicitly references it, but the discussion focuses on ski races, which is only loosely related to the skier’s role. The reference is valid but not fully accurate, so the score is 4-6. - Example 2: The image shows a tree with roots exposed, and the text references “the importance of deep roots” in discussing personal growth. While the image supports the metaphor, the connection is not fully fleshed out, leading to a score of 4-6. - Example 3: The image is of a sunset, and the text discusses the environmental impact of light pollution. While the sunset relates to the topic of light, it does not directly contribute to the explanation of light pollution, resulting in a score of 4-6. - Score 7-9: The image is explicitly referenced in a generally proper and correct manner. - Example 1: The image depicts a box of chips and is explicitly mentioned in the text at an appropriate point. While the reference is accurate, it might feel a little stiff, leading to a score of 7-9. - Example 2: The text discusses the structural components of a building, and the image shows a detailed diagram of the beams and supports. The reference is clear and mostly accurate, with a slight stiffness in phrasing, justifying a score of 7-9. - Example 3: The image depicts a city map, and the text discusses the layout of different districts. The map is explicitly referenced and properly aids the understanding of the text, but the reference could be smoother. The score should be 7-9. - Score 10: The image is explicitly referenced with complete accuracy and proper placement. - Example 1: The image is explicitly mentioned and thoroughly explained in the text, with a discussion of NFL play calls perfectly aligned with the visual content. The placement and explanation of the image are spot-on, so it warrants a score of 10. - Example 2: The image shows a detailed diagram of the human circulatory system, and the text provides an in-depth explanation of how blood flows through the heart and veins. The image and text work in perfect harmony to enhance the reader’s understanding. The score should be 10. - Example 3: The image depicts a vintage car and is referenced explicitly in the context, which discusses the history and evolution of automobile designs. The reference is fully accurate, and the image provides essential clarity to the reader, justifying a perfect score of 10. Note: The image only needs to be referenced or explained by its context above OR below, not necessarily both. Be rigorous and objective in your assessment. # Output Instructions
@@
E.2 Evaluation Costs
The average costs of our evaluation metrics are listed in Table 6.
Metric Cost (\$ per 1K samples)
Text-modal Metrics
Fluency 3.50
Relevance 3.58
Context Precision 4.61
Faithfulness 3.98
Multi-modal Metrics
Image Coherence 4.76
Image Helpfulness 4.92
Image Reference 4.70
Table 6: Evaluation costs of our metrics. The prices represent the average evaluation cost of the experiments using the joint modeling approach, as presented in Table  2. The text-modal metrics correspond to the evaluation cost of an entire generated response, whereas the multi-modal metrics correspond to the evaluation cost per image in the output. The Context Precision and Faithfulness metrics are based on the commercial model GPT-4o-mini, while all other metrics use GPT-4o.
Appendix F Experimental Setup